OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57 21 -21 9 -9

OpenStudy (anonymous):

help.... :=(

OpenStudy (anonymous):

can you help me ?

OpenStudy (anonymous):

sure, i will calculate here

OpenStudy (anonymous):

its negative right

OpenStudy (anonymous):

7x - 2y = 21 --> eq.[1] 4x + y = 57 --> eq.[2]

OpenStudy (anonymous):

from eq.[2] we have y = 57 - 4x. substitute it in eq.[1] we have: 7x -2(57 -4x) = 21 7x -144 +8x =21 15x = 135 x =9

OpenStudy (anonymous):

oh thanks i was using another meathod

OpenStudy (anonymous):

wc :)